OpenStudy (moongazer):

After taking off 20% from the original price of a pair of jeans, it cost PHP600. What is the original price?

OpenStudy (anonymous):

let original price be x takin off 20% would mean subtracting 20% of x from x i.e. 0.8x so 0.8x=600 x=750

OpenStudy (moongazer):

would you solve it easier? use fractions instead of decimals

OpenStudy (anonymous):

amount u paid was 80% of x then u can calculate the original price

OpenStudy (moongazer):

could you solve it in fraction form?

OpenStudy (anonymous):

ok x - (20/100)x=600 (80/100)x=600 x=600*(100/80) x=750 is it good?

OpenStudy (moongazer):

ahhhhh ok i was just confused ^_^ thanks!